you want to construct walkway 4 feet wide all the way around a pool. The rectangular shaped pool has a length that is 3 less than twice the width with a total area of 53 square feet. What are the dimensions of the pool? What are the dimensions of the walkway? what is the total area of concrete that will be constructed? Explain how you get every answer?
Answer by ankor@dixie-net.com(22740)

you want to construct walkway 4 feet wide all the way around a pool.
The rectangular shaped pool has a length that is 3 less than twice the width with a total area of 53 square feet.
What are the dimensions of the pool? What are the dimensions of the walkway?
what is the total area of concrete that will be constructed?
Explain how you get every answer?
:
Let L = the pool length
Let W = the pool width
"a length that is 3 less than twice the width"
L = 2W-3
"with a total area of 53 square feet."
L * W = 53
Replace L with (2W-3)
(2W-3) * W = 53
2W^2 - 3W = 53
2W^2 - 3W - 53 = 0
This won't factor; using the quadratic formula:
W = 5.952 ft is the width
Find the Length
L = 2(5.952)-3
L = 8.904 ft is the lenght
Check: 8.904*5.952 = 53.0
:
" What are the dimensions of the walkway?"
A 4'walkway all the way around will add 8' to the pool dimensions
16.904 by 13.952 ft is the outside dimensions of the walkway
:
"what is the total area of concrete that will be constructed?"
That would be the total area including the walkway - the pool area
16.904*13.952 - 53 = 182.8 sq/ft of concrete